OpenStudy (anonymous):

What is the average of all possible 5-digit number containing each of the digits 1,2,3,4, and 5?

OpenStudy (anonymous):

Will it just be 5^5 = 3125?

OpenStudy (welshfella):

the number of possible arrangements is 5! = 5*4*3*2*1 = 120

OpenStudy (anonymous):

How about if only the digit 5 can be repeated any number of times?